# Bondi mass, memory effect and balance law of polyhomogeneous spacetime

### v2

@@ 6. Gravitational wave memory and balance law
Since $\Psi_{4}^{1}$ is closely related to geodesic deviation equation, from the viewpoint of the detection of gravitational waves, $\Psi_{4}^{1}$ or $\sigma_{2,0}$ is the most significant quantity [43]. In fact, $\sigma_{2,0}$ is related to the plus mode $h_{+}$ and the cross mode $h_{\times}$ of gravitational waves by
@@
where $D_{s}$ is the luminosity distance between the observer and the source [27].
@@
ð$\eth^{2}\big{(}\bar{\sigma}_{2,0}\big{|}_{-\infty}^{\infty}\big{)}=-\bigg{(}\Psi_{2}^{3,0}+\sigma_{2,0}\dot{\bar{\sigma}}_{2,0}\bigg{)}\bigg{|}_{-\infty}^{\infty}+\int_{-\infty}^{\infty}|\dot{\sigma}_{2,0}|^{2}\mathrm{d}u.$ (6.5)
This formula describes the gravitational waves memory in polyhomogeneous spacetimes. It turns out that (6.4) and (6.5) remain unchanged as the ones obtained in smooth asymptotically flat spacetimes [27] and [24].
Pasterski, Strominger and Zhiboedov proposed a new type of gravitational memory named the spin memory [44]. It was shown that the effect of the spin memory is the relative time delay of different light rays at very large radial distance $r_{0}$. Mao and Wu found a new observational effect by looking at timelike geodesics near null infinity [45]. In the rest of this section, we extend Mao and Wu’s results to polyhomogeneous asymptotically flat spacetimes.
It can be obatined from Section 2 that the covariant metric components of a polyhomogeneous asymptotically flat spacetime are
$g_{00}=1-\frac{2U_{1}}{r}+\cdots,$ (6.6)
$g_{01}=1,$ (6.7)
$g_{02}=\bigg{(}X_{2}^{\theta}-\frac{\omega_{1}+\bar{\omega}_{1}}{\sqrt{2}}\bigg{)}+\cdots,$ (6.8)
$g_{03}=\bigg{(}X_{2}^{\varphi}\sin^{2}\theta+\frac{i(\omega_{1}-\bar{\omega})\sin\theta}{\sqrt{2}}\bigg{)}+\cdots,$ (6.9)
$g_{11}=0,g_{12}=0,g_{13}=0,$ (6.10)
$g_{22}=-r^{2}+\sqrt{2}(\xi_{2}^{\theta}+\bar{\xi}_{2}^{\theta})r+\cdots,$ (6.11)
$g_{23}=\frac{r\sin\theta}{\sqrt{2}}\bigg{(}i\bar{\xi}^{\theta}_{2}-i\xi_{2}^{\theta}+(\xi_{2}^{\varphi}+\bar{\xi}_{2}^{\varphi})\sin\theta\bigg{)}+\cdots,$ (6.12)
$g_{33}=-r^{2}\sin^{2}\theta-i\sqrt{2}r(\xi_{2}^{\varphi}-\bar{\xi}_{2}^{\varphi})\sin^{3}\theta+\cdots.$ (6.13)
Consider a timelike hypersurface $\Sigma_{r_{0}}=\{r=r_{0}\}$ in a polyhomogeneous spacetime with $r_{0}$ a large constant. Up to relevant orders, the induced metric $h_{ij}$ reads
$\mathrm{d}s_{\Sigma}^{2}$ $=$ $h_{ij}\mathrm{d}x^{i}\mathrm{d}x^{j}$ (6.14)
$=$ $\bigg{[}1-\frac{2U_{1}}{r}+\frac{1}{r^{2}}\bigg{(}-2U_{2}-(X_{2}^{\theta})^{2}+\sqrt{2}X_{2}^{\theta}(\omega_{1}+\bar{\omega}_{1})$
@@
$+2\bigg{[}\bigg{(}X_{2}^{\theta}-\frac{\omega_{1}+\bar{\omega}_{1}}{\sqrt{2}}\bigg{)}+\frac{1}{2r}\bigg{(}2X_{3}^{\theta}-2\sqrt{2}X_{2}^{\theta}(\xi_{2}^{\theta}+\bar{\xi}_{2}^{\theta})$
@@
$+2\bigg{[}\bigg{(}X_{2}^{\varphi}\sin^{2}\theta+\frac{i\sin\theta}{\sqrt{2}}(\omega_{1}-\bar{\omega}_{1})\bigg{)}-\frac{i\sin\theta}{2r}\bigg{(}2iX_{3}^{\varphi}\sin\theta$
@@
$+\bigg{[}-r^{2}+\sqrt{2}(\xi_{2}^{\theta}+\bar{\xi}_{2}^{\theta})r+\frac{1}{2}\bigg{(}-3(\xi_{2}^{\theta}+\bar{\xi}_{2}^{\theta})^{2}+2\sqrt{2}(\xi_{3}^{\theta}+\bar{\xi}_{3}^{\theta})$
@@
$+\bigg{[}\frac{r\sin\theta}{\sqrt{2}}\bigg{(}(\xi_{2}^{\varphi}+\bar{\xi}_{2}^{\varphi})\sin\theta-i(\xi_{2}^{\theta}-\bar{\xi}_{2}^{\theta})\bigg{)}-\frac{i\sin\theta}{2}\bigg{(}-(\xi_{2}^{\varphi})^{2}+(\bar{\xi}_{2}^{\varphi})^{2}$
@@
$+\bigg{[}-r^{2}\sin^{2}\theta-\sqrt{2}ir\sin^{3}\theta(\xi_{2}^{\varphi}-\bar{\xi}_{2}^{\varphi})+\frac{1}{2}\sin^{2}\theta\bigg{(}(\xi_{2}^{\theta}-\bar{\xi}_{2}^{\theta})^{2}$
@@
$+3(\xi_{2}^{\varphi}-\bar{\xi}_{2}^{\varphi})^{2}\sin^{2}\theta\bigg{)}+\cdots\bigg{]}\mathrm{d}\varphi^{2}.$
@@
where $D_{i}$ is the covariant derivative operator associated with the induce metric $h_{ij}.$ The asymptotic behaviour of $V^{a}$ are assumed as [45]
$V^{u}=1+\sum_{k=1}^{\infty}\frac{V_{k}^{u}(u,z,\theta,\varphi)}{r^{k}},$ (6.16)
$V^{\theta}=\sum_{k=1}^{\infty}\frac{V_{k}^{\theta}(u,z,\theta,\varphi)}{r^{k}},$ (6.17)
$V^{\phi}=\sum_{k=1}^{\infty}\frac{V_{k}^{\phi}(u,z,\theta,\varphi)}{r^{k}},$ (6.18)
where $z=\ln r.$
@@
$\bullet$ At order of $O(\frac{1}{r})$, we have
$\frac{\partial V_{1}^{\theta}(u,z,\theta,\varphi)}{\partial u}=0\ \Rightarrow\ V_{1}^{\theta}=0.$ (6.19)
$\frac{\partial V_{1}^{\varphi}(u,z,\theta,\varphi)}{\partial u}=0\ \Rightarrow\ V_{1}^{\varphi}=0.$ (6.20)
$\frac{\partial V_{1}^{u}(u,z,\theta,\varphi)}{\partial u}-\frac{\partial U_{1}(u,z,\theta,\varphi)}{\partial u}=0$
$\Rightarrow V_{1}^{u}(u,z,\theta,\varphi)=U_{1}(u,z,\theta,\varphi).$ (6.21)
@@
$\bullet$ At order of $O(\frac{1}{r^{2}})$, we have
@@
$\bullet$ At order of $O(\frac{1}{r^{3}})$, we have
@@
$\ \ \ \ \ \ \ \ \ \ \ -\xi_{2}^{\theta}\bar{\omega}_{1}-\bar{\xi}_{2}^{\theta}\omega_{1}+\int^{u}\frac{\partial U_{1}(v,z,\theta,\varphi)}{\partial\theta}\mathrm{d}v.$ (6.25)
@@
$\ \ \ \ \ \ \ \ \ \ \ \ -\xi_{2}^{\varphi}\bar{\omega}_{1}-\bar{\xi}_{2}^{\varphi}\omega_{1}+\frac{1}{\sin^{2}\theta}\int^{u}\frac{\partial U_{1}(v,z,\theta,\varphi)}{\partial\varphi}\mathrm{d}v.$ (6.26)
@@
$\mathcal{M}:=-\int^{u}U_{1}(v,z,\theta,\varphi)\mathrm{d}v.$ (6.28)
@@
This means the observer receives a time delay of order $O(\frac{1}{r})$. It should be emphasized that $\Delta\mathcal{M}$ is angle dependent.
@@
$\gamma_{2,1}=-\frac{1}{2}\eth^{\prime}\eth^{\prime}\Psi_{0}^{3,0}+\frac{1}{2}\dot{\bar{\sigma}}_{2,0}\Psi_{0}^{3,0},$ (6.33)
@@
ðð$\eth^{\prime}\eth\Delta\mathcal{M}$ $=$ $\Delta\bigg{[}\frac{1}{2}\eth^{\prime}\Psi_{1}^{4,0}+\frac{1}{2}\eth\bar{\Psi}_{1}^{4,0}+(\eth^{\prime}\sigma_{2,0})(\eth\bar{\sigma}_{2,0})\bigg{]}$ (6.36)
$+$ $\int_{u_{i}}^{u_{f}}\bigg{[}\sigma_{2,0}(\eth^{\prime}\eth\dot{\bar{\sigma}}_{2,0})+\bar{\sigma}_{2,0}(\eth\eth^{\prime}\dot{{\sigma}}_{2,0})\bigg{]}\mathrm{d}v$
$+$ $\int_{u_{i}}^{u_{f}}\bigg{[}\frac{1}{2}\eth^{\prime}(\Psi_{0}^{3,0}\eth\dot{\bar{\sigma}}_{2,0})+\frac{1}{2}\eth(\bar{\Psi}_{0}^{3,0}\eth^{\prime}\dot{{\sigma}}_{2,0})+\frac{3}{2}\eth^{\prime}\eth(\gamma_{2,1}+\bar{\gamma}_{2,1})\bigg{]}\mathrm{d}v$
$+$ $\bigg{[}\int_{u_{i}}^{u_{f}}\eth^{\prime}\eth(\gamma_{2,1}+\bar{\gamma}_{2,1})\mathrm{d}v\bigg{]}z.$
@@
ðð$\eth^{\prime}\eth\Delta\mathcal{M}$ $=$ $\Delta\bigg{[}\frac{1}{2}\eth^{\prime}\Psi_{1}^{4,0}+\frac{1}{2}\eth\bar{\Psi}_{1}^{4,0}+(\eth^{\prime}\sigma_{2,0})(\eth\bar{\sigma}_{2,0})\bigg{]}$ (6.37)
$+$ $\int_{u_{i}}^{u_{f}}\bigg{[}\sigma_{2,0}(\eth^{\prime}\eth\dot{\bar{\sigma}}_{2,0})+\bar{\sigma}_{2,0}(\eth\eth^{\prime}\dot{{\sigma}}_{2,0})\bigg{]}\mathrm{d}v.$
This is exactly the same as Eq.(4.8) in [45].
7. Conclusions
Spacetimes whose metrics admit an expansion in terms of a combination of powers of $1/r$ and $\ln r$ are called polyhomogeneous spacetimes. The asymptotic behaviour of the Newman-Penrose quantities, including the metric coefficients, the spin coefficients, and the tetrad components of the Weyl tensor are obtained in vacuum. The Bondi mass of the polyhomogeneous spacetimes is derived via the Iyer-Wald formalism. It is just the charge associated with the asymptotic translation at null infinity. It should be remarked that our method here differs from the previous work in [20, 21]. We deal with the physical spacetime and no conformal compactification is required. And the gauge choice we adopt is slightly different from the one used in [15]. Gravitational wave memory of the polyhomogeneous spacetimes are discussed. The appearance of the logarithmic terms in the expansion does not change the balance law.

### v3

@@ 6 Gravitational wave memory and balance law
Since $\Psi_{4}^{1}$ is closely related to geodesic deviation equation, from the viewpoint of the detection of gravitational waves, $\Psi_{4}^{1}$ or $\sigma_{2,0}$ is the most significant quantity Mag08 . In fact, $\sigma_{2,0}$ is related to the plus mode $h_{+}$ and the cross mode $h_{\times}$ of gravitational waves by
@@
where $D_{s}$ is the luminosity distance between the observer and the source Cao21-1 .
@@
$\eth^{2}\big(\bar{\sigma}_{2,0}\big|_{-\infty}^{\infty}\big)=-\bigg(\Psi_{2}^{3,0}+\sigma_{2,0}\dot{\bar{\sigma}}_{2,0}\bigg)\bigg|_{-\infty}^{\infty}+\int_{-\infty}^{\infty}|\dot{\sigma}_{2,0}|^{2}\mathrm{d}u.$ (6.5)
This formula describes the gravitational waves memory in polyhomogeneous spacetimes. It turns out that (6.4) and (6.5) remain unchanged as the ones obtained in smooth asymptotically flat spacetimes Cao21-1 and Frau92 .
Pasterski, Strominger and Zhiboedov proposed a new type of gravitational wave memory named the spin memory Pasterski2016 . It was shown that the effect of the spin memory is the relative time delay of different light rays at very large radial distance $r_{0}$. Mao and Wu found a new observational effect by looking at timelike geodesics near null infinity Mao-Wu-2019 . In the rest of this section, we extend Mao and Wu’s results to polyhomogeneous asymptotically flat spacetimes.
From Section 2, we can obtain the covariant metric components of a polyhomogeneous asymptotically flat spacetime. Now we consider a timelike hypersurface $\Sigma_{r_{0}}=\{r=r_{0}\}$ in a polyhomogeneous spacetime with $r_{0}$ a large constant. Up to relevant orders, the induced metric $h_{ij}$ reads
$\mathrm{d}s_{\Sigma}^{2}$ $=$ $h_{ij}\mathrm{d}x^{i}\mathrm{d}x^{j}$
$=$ $\bigg[1-\frac{2U_{1}}{r_{0}}+\frac{1}{r_{0}^{2}}\bigg(-2U_{2}-(X_{2}^{\theta})^{2}+\sqrt{2}X_{2}^{\theta}(\omega_{1}+\bar{\omega}_{1})$
@@
$+2\bigg[\bigg(X_{2}^{\theta}-\frac{\omega_{1}+\bar{\omega}_{1}}{\sqrt{2}}\bigg)+\frac{1}{2r_{0}}\bigg(2X_{3}^{\theta}-2\sqrt{2}X_{2}^{\theta}(\xi_{2}^{\theta}+\bar{\xi}_{2}^{\theta})$
@@
$+2\bigg[\bigg(X_{2}^{\varphi}\sin^{2}\theta+\frac{i\sin\theta}{\sqrt{2}}(\omega_{1}-\bar{\omega}_{1})\bigg)-\frac{i\sin\theta}{2r_{0}}\bigg(2iX_{3}^{\varphi}\sin\theta$
@@
$+\bigg[-r_{0}^{2}+\sqrt{2}(\xi_{2}^{\theta}+\bar{\xi}_{2}^{\theta})r_{0}+\frac{1}{2}\bigg(-3(\xi_{2}^{\theta}+\bar{\xi}_{2}^{\theta})^{2}+2\sqrt{2}(\xi_{3}^{\theta}+\bar{\xi}_{3}^{\theta})$
@@
$+\bigg[\frac{r_{0}\sin\theta}{\sqrt{2}}\bigg((\xi_{2}^{\varphi}+\bar{\xi}_{2}^{\varphi})\sin\theta-i(\xi_{2}^{\theta}-\bar{\xi}_{2}^{\theta})\bigg)-\frac{i\sin\theta}{2}\bigg(-(\xi_{2}^{\varphi})^{2}+(\bar{\xi}_{2}^{\varphi})^{2}$
@@
$+\bigg[-r_{0}^{2}\sin^{2}\theta-\sqrt{2}ir_{0}\sin^{3}\theta(\xi_{2}^{\varphi}-\bar{\xi}_{2}^{\varphi})+\frac{1}{2}\sin^{2}\theta\bigg((\xi_{2}^{\theta}-\bar{\xi}_{2}^{\theta})^{2}$
@@
$+3(\xi_{2}^{\varphi}-\bar{\xi}_{2}^{\varphi})^{2}\sin^{2}\theta\bigg)+\cdots\bigg]\mathrm{d}\varphi^{2}.$ (6.6)
@@
where $D_{i}$ is the covariant derivative operator associated with the induced metric $h_{ij}.$ The asymptotic behaviour of $V^{a}$ are assumed as Mao-Wu-2019
$V^{u}=1+\sum_{k=1}^{\infty}\frac{V_{k}^{u}(u,z_{0},\theta,\varphi)}{r_{0}^{k}},$ (6.8)
$V^{\theta}=\sum_{k=1}^{\infty}\frac{V_{k}^{\theta}(u,z_{0},\theta,\varphi)}{r_{0}^{k}},$ (6.9)
$V^{\phi}=\sum_{k=1}^{\infty}\frac{V_{k}^{\phi}(u,z_{0},\theta,\varphi)}{r_{0}^{k}},$ (6.10)
where $z_{0}=\ln r_{0}.$
@@
At order of $O(\frac{1}{r_{0}})$, we have
$\frac{\partial V_{1}^{\theta}(u,z_{0},\theta,\varphi)}{\partial u}=0\ \Rightarrow\ V_{1}^{\theta}=0.$ (6.11)
$\frac{\partial V_{1}^{\varphi}(u,z_{0},\theta,\varphi)}{\partial u}=0\ \Rightarrow\ V_{1}^{\varphi}=0.$ (6.12)
$\frac{\partial V_{1}^{u}(u,z_{0},\theta,\varphi)}{\partial u}-\frac{\partial U_{1}(u,z_{0},\theta,\varphi)}{\partial u}=0$
$\Rightarrow V_{1}^{u}(u,z_{0},\theta,\varphi)=U_{1}(u,z_{0},\theta,\varphi).$ (6.13)
@@
At order of $O(\frac{1}{r_{0}^{2}})$, we have
@@
At order of $O(\frac{1}{r_{0}^{3}})$, we have
@@
$\ \ \ \ \ \ \ \ \ \ \ -\xi_{2}^{\theta}\bar{\omega}_{1}-\bar{\xi}_{2}^{\theta}\omega_{1}+\int^{u}\frac{\partial U_{1}(v,z_{0},\theta,\varphi)}{\partial\theta}\mathrm{d}v.$ (6.17)
@@
$\ \ \ \ \ \ \ \ \ \ \ \ -\xi_{2}^{\varphi}\bar{\omega}_{1}-\bar{\xi}_{2}^{\varphi}\omega_{1}+\frac{1}{\sin^{2}\theta}\int^{u}\frac{\partial U_{1}(v,z_{0},\theta,\varphi)}{\partial\varphi}\mathrm{d}v.$ (6.18)
@@
$\mathcal{M}:=-\int^{u}U_{1}(v,z_{0},\theta,\varphi)\mathrm{d}v.$ (6.20)
@@
This means the observer receives a time delay of order $O(\frac{1}{r_{0}})$. It should be emphasized that $\Delta\mathcal{M}$ is angle dependent.
@@
$\gamma_{2,1}=-\frac{1}{2}\bar{\eth}\bar{\eth}\Psi_{0}^{3,0}+\frac{1}{2}\dot{\bar{\sigma}}_{2,0}\Psi_{0}^{3,0},$ (6.25)
@@
$\bar{\eth}\eth\Delta\mathcal{M}$ $=$ $\Delta\bigg[\frac{1}{2}\bar{\eth}\Psi_{1}^{4,0}+\frac{1}{2}\eth\bar{\Psi}_{1}^{4,0}+(\bar{\eth}\sigma_{2,0})(\eth\bar{\sigma}_{2,0})\bigg]$ (6.28)
$+$ $\int_{u_{i}}^{u_{f}}\bigg[\sigma_{2,0}(\bar{\eth}\eth\dot{\bar{\sigma}}_{2,0})+\bar{\sigma}_{2,0}(\eth\bar{\eth}\dot{{\sigma}}_{2,0})\bigg]\mathrm{d}v$
$+$ $\int_{u_{i}}^{u_{f}}\bigg[\frac{1}{2}\bar{\eth}(\Psi_{0}^{3,0}\eth\dot{\bar{\sigma}}_{2,0})+\frac{1}{2}\eth(\bar{\Psi}_{0}^{3,0}\bar{\eth}\dot{{\sigma}}_{2,0})+\frac{3}{2}\bar{\eth}\eth(\gamma_{2,1}+\bar{\gamma}_{2,1})\bigg]\mathrm{d}v$
$+$ $\bigg[\int_{u_{i}}^{u_{f}}\bar{\eth}\eth(\gamma_{2,1}+\bar{\gamma}_{2,1})\mathrm{d}v\bigg]z_{0}.$
@@
$\bar{\eth}\eth\Delta\mathcal{M}$ $=$ $\Delta\bigg[\frac{1}{2}\bar{\eth}\Psi_{1}^{4,0}+\frac{1}{2}\eth\bar{\Psi}_{1}^{4,0}+(\bar{\eth}\sigma_{2,0})(\eth\bar{\sigma}_{2,0})\bigg]$ (6.29)
$+$ $\int_{u_{i}}^{u_{f}}\bigg[\sigma_{2,0}(\bar{\eth}\eth\dot{\bar{\sigma}}_{2,0})+\bar{\sigma}_{2,0}(\eth\bar{\eth}\dot{{\sigma}}_{2,0})\bigg]\mathrm{d}v.$
This is exactly the same as Eq.(4.8) in Mao-Wu-2019 .
As shown in Mao-Wu-2019 , by choosing a ring of freely falling observers who are initially static and synchronised, the difference in the proper time between each observer on this ring at a later time encodes the gravitational wave-form, providing an observable signature of spin memory.
7 Conclusions
In this article, we investigate the mass and the memory effect of polyhomogeneous spacetimes. The gravitational waves persist all the way to the null infinity. The analysis of the solutions of the Einstein field equations was studied in the basic settings by Bondi et al. Bondi62 ; Sachs62 ; NP62 using the formal series expansions in powers of $1/r$. Carving out the feature of null geometry, Penrose P63 proposed the idea of conformal compactifications. It requires that the conformal structure should be smoothly extended aross the null infinity. The implicit assumption of the smoothness of the coformal extension leads to the very specific fall-off behaviours of isolated gravitating bodies and the emitting radiation. This is knowns as the peeling-off property. If there is a slower decay assumption for the component of the Weyl tensors, then the logarithmic terms will appear in the asymptotic expansions. It seems that, regardless of the geometric elegance, the conformal compactification is a guess proposal and the purely $1/r$ expansions are too stringent. One kind of spacetimes removing the smoothness assumption are those that admit a polyhomogeneous expansion. Precisely, the metrics can be expanded in terms of a combination of powers of $1/r$ and $\ln r$.
The asymptotic behaviour of the Newman-Penrose quantities of polyhomogeneous spacetime, including the metric coefficients, the spin coefficients, and the tetrad components of the Weyl tensor are obtained in vacuum. This generalization is nontrivial since the calculational complexity substantially increases when considering polyhomogeneous expansions. For smooth asympotically flat spacetimes, we have reexamined the Bondi mass at null infinity via the Iyer-Wald formula within the Bondi-Sachs coordinates. Our analysis demonstrates that the Iyer-Wald formula naturally leads to the Bondi mass expression directly in the physical spacetime, avoiding the difficulties and subtle isuues arising from the (unphysical) conformal compactification Grant2022 .
For polyhomogeneous spacetimes, we calculate the charge associated with the asymptotic translation at null infinity via the Iyer-Wald formalism and obtain the canonical energy $\mathcal{E}$ within the Newman-Unti gauge. The canonical energy $\mathcal{E}$ reduces to the Newman-Unti mass when the logarithmic terms vanish. However, the canonical energy does not has the monotonically decreasing property with respect to the retraded time. To find an energy expression which admits a mass-loss formula, we propose a new notion of energy $\hat{\mathcal{E}}$, called the ‘reduced energy’. This gives a possible way to define the Bondi mass of polyhomogeneous spacetimes. In particular, for polyhomogeneous spacetimes with $\Psi_{0}^{3}=0$, the result is consistent with the one obtained by Godazgar and Macaulay in (Godazagar2020, , Eq.(5.10)) using the Barnich-Brandt prescription Barnich2002 . It should also be remarked that our method here differs from the previous work in Wald2000 ; Grant2022 and the gauge choice we adopt is slightly different from the one used in Kroon99 .
Gravitational wave memory of polyhomogeneous spacetimes are discussed. In conclusion, we find the appearance of the logarithmic terms in the expansion does not change the balance law. This is a mathematical consequence of the symmetries of the Bianchi identities. More precisely, this result relies on the evolution equation for $\Psi_{2}^{3}+\sigma_{2}\bar{\sigma}_{2}$ which involves only the leading-order shear $\sigma_{2}(z)$ and its derivatives. The logarithmic terms in the full expansion of $\sigma_{2}(z)$ are of higher order in $1/r$ and do not have contributions in the memory effect. It indicates that the gravitational memory seems a good physically measurable effect even for spacetimes with slower decays.